OpenStudy (anonymous):

No matter how hard I try I can not figure out how to factor this: 4a^2-32a+63. Any help please? Thanks so much. Please explain the steps!

OpenStudy (anonymous):

(2 a-9) (2 a-7)

OpenStudy (anonymous):

So take the coefficient on the a squared term and multiply that to 63 you need to find two number (we will call them A and B) such that A * B = 4 * 63 and A + B = -32

OpenStudy (anonymous):

you do this with other polynomials but the coefficient on the squared term is usually a 1so most people don't teach it that way.

OpenStudy (anonymous):

You are fabulous. Thank you so much! No matter how hard I tried, I could not understand it. You've made it so much simpler!

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

I do have one more question...The two numbers I got were 14 and 18.. How did you get 7 and 9 from that?

OpenStudy (precal):

znimon is busy at the moment, have you tried to do factor by grouping

OpenStudy (precal):



OpenStudy (precal):